# Completed Feature Disentanglement Learning for Multimodal MRIs Analysis

### v1

@@ 4.2.3 Compared Methods
We compare the proposed framework with seven state-of-the-art (SOTA) MML methods, including EmbraceNet [20], ETMC [16], ADCCA [21], MAML [17], NestedFormer [18], MISA [8], DMD [10]. Specifically, ETMC is a uncertainty-based MML method. ADCCA is a CCA-based MML method. MAML and NestedFormer are attention-based MML methods originally designed for segmentation, but adapted to the classification task by adding a classifier after the encoders. MISA and DMD are FD-based MML methods. To ensure a fair comparison, in addition to the transformer-based NestedFormer method, we set the backbones of other CNN-based comparison methods to be the same as that of the proposed framework.
@@
Figure 4: Visualization results of the proposed framework on the MRNet and MEN datasets. In each sub-figure, the left part displays t-SNE visualization of the decoupled features, and the right part shows the heatmap which displays the cosine similarity of each pair-wise decoupled features. In the heatmap, the closer the color is to yellow, the greater the similarity between the two features. Conversely, the closer the color is to blue, the smaller the similarity between the two features.
@@
Table 2: The comparison results on the MEN dataset (mean±$\pm$±standard deviation). The best and second best results for each metric are respectively highlighted by red and blue.
@@ 4.3.1 Evaluation on Meniscal Tear Prediction Dataset
The comparison results on the MRNet dataset are summarized in Table 1. Among the comparison methods, the FD-based method DMD [10] achieves the best AUC ($0.7668$), while the uncertainty-based method ETMC [16] obtains the second place ($0.7655$). The attention-based method MAML [17] obtains better results than other comparison methods in three metrics: ACC ($0.7306$), Ba_ACC ($0.7261$) and AUPRC ($0.6148$). Our proposed framework achieves the first place in five metrics: ACC ($0.7389$, $0.0083$ better than the 2nd), G-Mean ($0.7351$, $0.0096$ better than the 2nd), Ba_ACC ($0.7372$, $0.0111$ better than the 2nd), AUPRC ($0.6207$, $0.0059$ better than the 2nd) and AUC ($0.8029$, $0.0361$ better than the 2nd). The higher ACC and AUC, along with the more balanced accuracy between positive and negative cases, demonstrate the effectiveness of our proposed framework on the MRNet dataset.
@@ 4.4 Ablation Analysis
Table 3: The ablation study of the proposed framework on the MRNet dataset.
@@
Table 4: The ablation study of the proposed framework on the MEN dataset.
@@
Figure 5: Heatmaps of feature weights in DMF module on the MRNet and MEN datasets. (a) and (c) show the mean feature weights of all test data on both datasets. (b) and (d) display the feature weights of four cases in each dataset, respectively.
@@ 4.4.1 Effectiveness of the CFD Strategy
Furthermore, we conduct ablation studies for $\alpha$ and $\beta$ on both adopted datasets, as shown in Fig. 4. Specifically, $\alpha$ and $\beta$ are balance factors of losses related to the CFD strategy ($\mathcal{L}_{sh}$, $\mathcal{L}_{ps}$ and $\mathcal{L}_{diff}$). For both datasets, the comparison results (see sub-figure (a) and (b) for MRNet dataset and sub-figure (c) and (d) for MEN dataset) clearly demonstrate that both modality-shared features and modal-partial-features are better learnt when using CFD-related losses ($\alpha\neq 0$, $\beta\neq 0$).
@@ 4.4.2 Effectiveness of the DMF Module
Moreover, we draw heatmaps of weights for the final decoupled features learned in the LinG_GN. Fig. 5 (a) and (b) show the heatmaps on the MRNet dataset. Firstly, we plot the mean weight of all cases in the test set for each final decoupled feature in a heatmap named MRNet-mean (see Fig. 5 (a)). This heatmap illustrates that $G_{12}$ and $P_{2}$ play greater roles during feature integration, with $G_{12}$ obtaining a maximum weight of around $0.4$. Specifically, $G_{12}$ represents the final modality-partial-shared feature between PD-axial and T1-corona, and $P_{2}$ represents modality-specific feature of T1-corona. Additionally, we randomly display the heatmaps of four cases (see Fig. 5 (b)).
@@
These heatmaps on both adopted datasets demonstrate that our proposed DMF module can dynamically capture the relationships between the decoupled features across different samples (see Fig. 5 (b) for the MRNet dataset and Fig. 5 (d) for the MEN dataset). Moreover, there are modality-partial-shared features playing important roles during feature fusion on both datasets, further illustrating the necessity of the CFD strategy.
@@ 5.1 Evaluation in the Four-modal Case
However, it is important to acknowledge that as the number of modalities increases, the relationships between multiple modalities become more intricate, posing challenges for the CFD strategy. The inclusion of more modalities as inputs leads to an increase in network parameters, thereby exacerbating the difficulty of network optimization. However, in clinical studies, commonly used multimodal MRI datasets typically contain two to four modalities. Our proposed framework has been designed to achieve effective performance under these conditions.
@@
Table 5: The comparison results on the BraTS 2021 dataset (mean±$\pm$±standard deviation). The best and second best results for each metric are highlighted by red and blue.
@@ 5.2 Computational Complexity Analysis
Table 6: The computational complexity comparison between proposed and comparison methods in the three-modal case and four-modal case.
Methods Three-modal Four-modal
Parameters (M) GFLOPS Parameters (M) GFLOPS
EmbraceNet [20] 99.62 46.16 132.78 32.23
MISA [8] 100.72 46.17 134.22 45.44
MAML [17] 103.26 165.63 137.63 111.93
ETMC [16] 101.62 46.17 135.44 45.44
ADCCA [21] 99.93 46.16 133.24 45.44
DMD [10] 104.22 46.27 142.75 45.67
Proposed 99.62 46.16 132.89 45.44
As shown in Table 6, we compare the number of parameters and GFLOPS between the proposed method and the comparison methods. Specifically, we only consider the comparison methods that adopt the same backbone as the proposed method. In the three-modal case, the proposed method achieves best performance with the fewest parameters and the lowest GFLOPS. In the four-modal case, the EmbraceNet method has the fewest parameters and GFLOPS, but our proposed method achieves a significant performance improvement using a similar number of parameters and the second-fewest GFLOPS. The number of parameters of all methods increases from the three-modal case to the four-modal case. However, our proposed method achieves a better performance improvement with a relatively small increase in parameters.
@@ 6 Conclusion and Future Work
In this paper, we propose an effective MML framework named CFDL. More importantly, we present a novel CFD strategy that separates multimodal information into modality-shared features, modality-specific features and modality-partial-shared features, where the latter are overlooked in previous FD-based methods. Our analysis and experiments demonstrate the significant roles played by modality-partial-shared features in prediction. Additionally, we present the DMF module, which dynamically fuses decoupled features in an explicit manner. The proposed LinG_GN in the DMF module can generate weights for the decoupled features by learning local-global relationship between these features. This customized fusion module can provide interpretability for clinical analysis, enabling a deeper understanding of the characteristics and behaviors of each decoupled features. Experimental results on three multimodal MRI datasets verify the effectiveness of the proposed framework. Moreover, the visualization results provide evidence for the benefits of our CFD strategy in achieving comprehensive and informative feature decoupling, and the DMF module in enabling dynamic fusion of the decoupled features. Furthermore, we consider that the underlying principles of proposed framework can be extended to other medical imaging tasks. In the future, we plan to explore the application of our framework to medical segmentation tasks, which are closely related to medical classification tasks.

### v2

@@ 4.2.3 Compared Methods
We compare the proposed framework with nine state-of-the-art (SOTA) MML methods, including EmbraceNet [21], ETMC [16], ADCCA [22], MAML [18], NestedFormer [19], MISA [8], DMD [10], CCML [17] and GLoMo [23]. Specifically, ETMC and CCML are uncertainty-based MML methods. ADCCA and GLoMo are correlation-based MML methods. MAML and NestedFormer are attention-based MML methods originally designed for segmentation, but adapted to the classification task by adding a classifier after the encoders. MISA and DMD are FD-based MML methods. To ensure a fair comparison, in addition to the transformer-based NestedFormer method, we set the backbones of other CNN-based comparison methods to be the same as that of the proposed framework.
@@
Figure 5: Visualization results of the proposed framework on the MRNet and MEN datasets. Each sub-figure includes t-SNE visualization of the decoupled features (left) and a heatmap of the cosine similarities for each pair-wise decoupled features (right). In the heatmap, yellow indicates higher similarity, while blue indicates lower similarity.
@@
Table 2: The comparison results on the MEN dataset.
@@ 4.3.1 Evaluation on Meniscal Tear Prediction Dataset
The comparison results on the MRNet dataset are summarized in Table 1. Among the comparison methods, the correlation-based method GLoMo [23] achieves the best AUC ($0.7670$), while the uncertainty-based method CCML [17] obtains the second place ($0.7669$). The attention-based method MAML [18] obtains better results than other comparison methods in three metrics: ACC ($0.7306$), Ba_ACC ($0.7261$) and AUPRC ($0.6148$). Our proposed framework achieves the first place in five metrics: ACC ($0.7389$, $0.0083$ better than the 2nd), G-Mean ($0.7351$, $0.0096$ better than the 2nd), Ba_ACC ($0.7372$, $0.0111$ better than the 2nd), AUPRC ($0.6207$, $0.0059$ better than the 2nd) and AUC ($0.8029$, $0.0359$ better than the 2nd). The higher ACC and AUC, along with the more balanced accuracy between positive and negative cases, demonstrate the effectiveness of proposed framework on the MRNet dataset.
@@ 4.4 Ablation Analysis
Table 3: The ablation study on the MRNet dataset.
@@
Table 4: The ablation study on the MEN dataset.
@@
Figure 6: Heatmaps of feature weights in DMF module on the MRNet and MEN datasets. (a) and (c) show the mean feature weights of all test data on both datasets. (b) and (d) display the feature weights of two cases in each dataset, respectively.
@@ 4.4.1 Effectiveness of the CFD Strategy
Furthermore, we conduct ablation studies for $\alpha$ and $\beta$ on both adopted datasets, as shown in Fig. 5. Specifically, $\alpha$ and $\beta$ are balance factors of losses related to the CFD strategy ($\mathcal{L}_{sim}$ and $\mathcal{L}_{diff}$). For both datasets, the comparison results (see sub-figure (a) and (b) for MRNet dataset and sub-figure (c) and (d) for MEN dataset) clearly demonstrate that both modality-shared features and modal-partial-features are better learnt when using CFD-related losses ($\alpha\neq 0$, $\beta\neq 0$).
@@ 4.4.2 Effectiveness of the DMF Module
Moreover, we draw heatmaps of weights for the final decoupled features learned in the LinG_GN. Fig. 6 (a) and (b) show the heatmaps on the MRNet dataset. Firstly, we plot the mean weight of all cases in the test set for each final decoupled feature in a heatmap named MRNet-mean (see Fig. 6 (a)). This heatmap illustrates that $G_{12}$ and $P_{2}$ play greater roles during feature integration, with $G_{12}$ obtaining a maximum weight of around $0.4$. Specifically, $G_{12}$ represents the final modality-partial-shared feature between PD-axial and T1-corona, and $P_{2}$ represents modality-specific feature of T1-corona. Additionally, we randomly display the heatmaps of two cases (see Fig. 6 (b)).
@@
These heatmaps demonstrate that our proposed DMF module can dynamically capture the relationships between the decoupled features across different samples (see Fig. 6 (b) for the MRNet dataset and Fig. 6 (d) for the MEN dataset). Moreover, there are modality-partial-shared features playing important roles during feature fusion on both datasets, further illustrating the necessity of the CFD strategy.
@@ 5 Discussion
Figure 7: Visualization of activation maps for MRI modalities on the MEN dataset using Grad-CAM. Original images are shown on the left, with corresponding Grad-CAM images on the right. Darker red regions in the Grad-CAM images indicate areas of highest contribution to the prediction. In the FLAIR-C images, tumor and edema regions are marked with red and green boxes, respectively.
5.1 Visualization analysis
To enhance the interpretability of our framework, Grad-CAM [44] was employed to visualize the activation maps of the network for each modality in the MEN dataset. As shown in Fig 7, the model pays more attention to the tumor-edema junction across different modalities, with variations in the specific areas of focus. For the T1C modality, the model predominantly focuses on the tumor and its surrounding region. In the FLAIR-C modality, the model highlights the tumor and edema regions. Conversely, in the ADC modality, the model primarily attends to the edema region. Moreover, based on the feature weights in the DMF module (Fig. 6 (c)), the characteristics of the tumor-edema junction ($F$, the final modality-shared feature), the edema ($G_{23}$, the final modality-partial-shared feature between FLAIR-C and ADC) and the tumor-surrounding region ($P1$, the modality-specific feature of T1C) contribute significantly to the final task.
Figure 8: Data distributions in intermediate layers. The term Bm-C2 denotes the output from the $2$nd convolutional layer of the $m$th block within the ResNet18 architecture of ADC modal branch. Each point on the graph corresponds to an individual sample, with points differentiated by color to indicate various categories.
Figure 9: Data distributions in the second last FC layer with different training epochs.
To validate the effectiveness of our framework, we visualized the feature space distribution using the Manifold Discovery and Analysis (MDA) [45] algorithm on the MEN dataset, randomly selecting balanced samples across categories for optimal visualization. Fig. 8 demonstrates that categorical separation becomes increasingly distinct in deeper network blocks, with B2-C2 through B5-C2 showing progressive improvement in class discrimination. The temporal evolution of feature distributions across training epochs (Fig. 9 (a-c)) reveals increasingly pronounced categorical boundaries. Furthermore, we examined the model’s robustness by analyzing feature distributions under Gaussian noise perturbation of inputs, as shown in Fig. 9 (d). The results showed that the feature distributions maintained clear categorical separation, demonstrating the method’s resilience to input variations.
@@ 5.2 Evaluation in the Four-modal Case
However, it is important to acknowledge that as the number of modalities increases, the relationships between them become more intricate, posing challenges for the CFD strategy. Additionally, incorporating more modalities increases the number of network parameters, further complicating network optimization. However, in clinical studies, commonly used multimodal MRI datasets typically contain two to four modalities. Our proposed framework has been designed to achieve effective performance under these conditions.
@@
Table 5: The comparison results on the BraTS 2021 dataset.
@@
Figure 11: The computational complexity comparison including parameter and GFLOPs between proposed and comparison methods in the three-modal case (left part) and four-modal case (right part).
@@ 5.3 Computational Complexity Analysis
As shown in Fig. 11, we compare the number of parameters and GFLOPs between the proposed method and the comparison methods [^4] We compare only with methods that utilize the same backbone as the proposed approach. For clarity, MAML is not shown in Fig. 11, as its parameter and GFLOPs values are significantly larger than those of other methods. Specifically, in the three-modality case, MAML has 202.61M parameters and 165.63 GFLOPs, while in the four-modality case, these values increase to 270.09M parameters and 167.89 GFLOPs.. In the three-modal case (left part of Fig. 11), using MRNet as an example, the proposed method achieves the best performance with the fewest parameters and the lowest GFLOPs. In the four-modal case (right part of Fig. 11), while EmbraceNet and CCML have the fewest parameters and GFLOPs, our method achieves a significant performance improvement with a comparable number of parameters and relatively low GFLOPs. Although the number of parameters increases for all methods from the three-modal case to the four-modal case, our proposed method achieves superior performance with only a modest increase in parameters.
@@ 6 Conclusion and Future Work
In this paper, we propose an effective MML framework called CFDL, incorporating a novel CFD strategy that separates multimodal information into modality-shared, modality-specific, and modality-partial-shared features, the last of which has been overlooked in previous FD-based methods. Our analysis and experiments demonstrate the critical role of modality-partial-shared features in prediction. Additionally, we present the DMF module, which explicitly and dynamically fuses the decoupled features. The LinG_GN within the DMF module can generate the decoupled feature weights by capturing their local-global relationships. This customized fusion module can provide interpretability for clinical analysis, enabling a deeper understanding of the characteristics and behaviors of each decoupled features. Furthermore, we consider that the underlying principles of proposed framework can be extended to other medical imaging tasks. In the future, we plan to explore the application of our framework to medical segmentation tasks, which are closely related to medical classification tasks.